A person invested $10,000. The greater part is invested at 6% and the remainder at 5%. The annual income from the greater part is $160 more than the annual income from the smaller part. What is the smaller part of the investement?
Answer by mananth(16949)

Part I 6.00% per annum------------- x
Part II 5.00% per annum------------ y
total investment 10000
Interest----- from greater part = 6.00%of x
Interest from smaller part= 5.00% of y
greater by $160
6.00%x= 5.00%y+160
multiply by 100
6x-5y= 16000 --------1
x+y=10000 -------- 2
eliminate y
Multiply 1 by 1
Multiply 2 by 5
we get
6x-5y=16000
5x-5y=50000
11x=66000
/11
x=6000
Plug the value of x in in 1
36000-5y=16000
-5y=16000-36000
-5y=-20000
/-5
y=4000
$6000 at 6.00%
$4000 at 5.00%